### v2

# OmniSVG: A Unified Scalable Vector Graphics Generation Model

Abstract
Scalable Vector Graphics (SVG) is an important image format widely adopted in graphic design because of their resolution independence and editability. The study of generating high-quality SVG has continuously drawn attention from both designers and researchers in the AIGC community. However, existing methods either produces unstructured outputs with huge computational cost or is limited to generating monochrome icons of over-simplified structures. To produce high-quality and complex SVG, we propose OmniSVG, a unified framework that leverages pre-trained Vision-Language Models (VLMs) for end-to-end multimodal SVG generation. By parameterizing SVG commands and coordinates into discrete tokens, OmniSVG decouples structural logic from low-level geometry for efficient training while maintaining the expressiveness of complex SVG structure. To further advance the development of SVG synthesis, we introduce MMSVG-2M, a multimodal dataset with two million richly annotated SVG assets, along with a standardized evaluation protocol for conditional SVG generation tasks. Extensive experiments show that OmniSVG outperforms existing methods and demonstrates its potential for integration into professional SVG design workflows.
Figure 1: Highlighted features of OmniSVG. OmniSVG is capable of autoregressively generating high-quality SVGs across a wide spectrum of complexity — from simple icons to intricate anime characters. It demonstrates remarkable versatility through multiple generation modalities, including Text-to-SVG, Image-to-SVG, and Character-Reference SVG, making it a powerful and flexible solution for diverse creative tasks.
footnotetext: * Yiying Yang and Wei Cheng contributed equally to this work.footnotetext: ‡ Corresponding Authors.
1 Introduction
Scalable Vector Graphics (SVG) have become a cornerstone of modern digital design because of their resolution independence, compact file size, and inherent editability. Widely adopted in professional workflows from UI/UX design to industrial CAD systems, SVG enables precise manipulation of geometric primitives (e.g., Bézier curves, polygons) while maintaining high precision and consistent visual quality across varying resolutions. However, creating high-quality SVG content remains challenging for non-experts, requiring mastery of specialized tools or intricate XML syntax.
Existing methods adopt either optimization-based methods or auto-regressive approaches to generate SVG contents. The optimization-based methods [21, 33, 59] iteratively refine the SVG parameters by optimizing differentiable vector graphics rasterizers. Though these methods are effective in generating SVG icons, they suffer from significant computational overhead when scaling up to intricate samples and produce unstructured outputs with redundant anchor points. In contrast, auto-regressive methods build transformer models or adapt pre-trained Large Language Models (LLMs) to directly generate XML parameters [58] or codes [56, 41] representing SVGs. Benefiting from the end-to-end learning pipeline, the auto-regressive method is a more scalable approach [5] as it can learn directly from a large collection of SVG samples. However, existing auto-regressive approaches are limited to basic SVG contents [11, 22, 52] because of the limited context length and the scarcity of complex SVG data.
In this paper, we propose OmniSVG that harnesses native VLMs [1] for various end-to-end multimodal SVG generation tasks. By parameterizing SVG coordinates and commands into discrete tokens, OmniSVG decouples structural logic from low-level geometry, mitigating the “coordinate hallucination” problem prevalent in code-based LLMs, and produces vivid and colorful SVG results. Additionally, the next token prediction training objective enables OmniSVG to complete SVGs with diverse generation results given some partial observations. Compared to traditional auto-regressive SVG generation methods, OmniSVG is able to parameterize SVGs exceeding $30k$ tokens, facilitating the generation of detailed and complex SVG contents. Building upon pre-trained VLMs, our method natively integrates the ability to reason upon visual and textual instructions to synthesize editable, high-fidelity SVGs across diverse domains, from icons to intricate illustrations and anime characters.
To advance the development of SVG synthesis, we introduce MMSVG-2M, a multi-modal SVG synthesis dataset with two million richly annotated assets, encompassing icons, illustrations, and anime designs. We also establish a standardized evaluation protocol, MMSVG-Bench, for three key SVG generation tasks, i.e. “Text-to-SVG”, “Image-to-SVG”, and “Character Reference SVG Generation”. Extensive experiments show that OmniSVG produces highly detailed and complex SVG contents, surpassing prior art both quantitatively and qualitatively.
To summarize, our key contributions include:
  • We introduce OmniSVG, a family of end-to-end multimodal SVG generators that leverage native VLMs for generating complex and detailed SVGs, from simple icons to intricate anime characters.
  • We present MMSVG-2M, a large-scale dataset comprising two million SVG assets, along with a standardized evaluation protocol for various multi-modal SVG generation tasks, providing a comprehensive resource for future research.
  • Extensive experiments show that OmniSVG surpasses prior SVG generation methods both qualitatively and quantitatively, highlighting its potential for integration into professional SVG design workflows.
2 Related Works
SVG Generation. Early attempts to generating SVGs directly utilize architectures like RNNs [17, 40, 18, 43, 44], VAEs [4, 31, 48, 45, 50, 65], and Transformers [4, 55] to compress SVG commands into latent representations. Meanwhile, DeepSVG [4] further parameterizes SVGs using a dual transformer architecture but struggles with geometric consistency. Recently, the advent of large language models (LLMs) [29, 63, 51, 60, 5, 6, 62, 61, 47] unleashes the potential of generating SVGs via XML code synthesis [58, 56, 41]. However, the limited context length of existing LLM-based SVG generation methods [56, 41, 58] poses significant challenges in handling complex SVGs that exceed $10k$ tokens. In this paper, we explore the potential of native Vision-Language Models (VLMs) in multi-modal SVG generation. Combining pre-trained VLMs with SVG command parameterization, we validate that OmniSVG is able to follow multi-modal instructions and generate more complex SVGs than before.
Image Vectorization. Recent advancements in vectorization harness diffusion models paired with differentiable rasterizers, using techniques like score distillation sampling [36, 21, 7] and specialized regularizers [27, 33] to convert raster images into SVG paths. While these methods achieve remarkable results, they face limitations such as over-smoothing, color over-saturation, and lack of editability, often producing tangled paths that fail to capture hierarchical structures inherent in professional SVG designs. In this paper, we present an end-to-end generation method that directly follows multi-modal instructions to generate high-quality SVGs with improved path clarity.
SVG Datasets and Benchmarks. The lack of suitable datasets for complex SVG structures presents a significant challenge. Existing datasets [11, 22, 52] primarily focus on simplified path-based SVGs or monochrome icons, overlooking the intricate layered structures and rich color semantics found in real-world designs. For example, FIGR-8-SVG [11] focuses on monochromatic icons, while StarVector [41] proposes categorized datasets, including illustrations, icons, emojis, and fonts. Therefore, existing datasets only present simple SVG samples that do not exceed $8.2k$ tokens, failing to capture the complexities of layered structures and rich color semantics. Benchmark evaluations, such as VGBench [71], further highlight gaps in multi-format testing and the absence of comprehensive coverage for illustrative SVGs. To this end, we introduce MMSVG-2M, a multimodal SVG synthesis dataset comprising two million richly annotated assets, including icons, illustrations, and complex anime designs. We also present a standardized evaluation protocol, MMSVG-Bench, to evaluate diverse multi-modal SVG generation tasks with varying complexity.
Figure 2: Overview of OmniSVG. OmniSVG is built on a pre-trained vision-language model Qwen2.5-VL and incorporates an SVG tokenizer. The model tokenizes both text and image inputs as prefix tokens, while the SVG tokenizer encodes vector graphics commands into a unified representation space.
3 OmniSVG Dataset
MMSVG-2M is a large-scale SVG dataset with two million SVG samples covering website icons, illustrations, graphic designs, anime characters, and so on (Sec. 3.1). To further promote the downstream development of SVG generation methods, we also introduce MMSVG-Bench, a benchmark consisting of a series of multi-modal instruction following tasks for conditional SVG generation (Sec. 3.2).
3.1 MMSVG-2M
Data Source.
The raw SVG we collected from the Internet consists of three types: MMSVG-Icon sourced from Iconfont, SVG illustrations sourced from iconsount, and SVG anime characters sourced from Freepik and our data creation pipeline depicted in Fig. 2. All these websites are online platforms where users can publish and share SVGs, encompassing a broad variety of categories. Specifically, our dataset comprises 2 million SVG samples within various subsets, including $1.1$ million SVG icons, $0.5$ million SVG illustrations, and 0.4 million SVG anime characters. We provide the data statistics of MMSVG-2M in Tab. 1.
Table 1: Data statistics for MMSVG-2M. Our MMSVG-2M consists of 1.1 million SVG icons, 0.5 million SVG illustrations, and 0.4 million SVG anime characters.
Dataset Train Val Test Source Token Length
MMSVG-Icon 990k 106.7k 3.3k Iconfonthttps://www.iconfont.cn/ 2.2k $\pm$ 0.9k
MMSVG-Illustration 450k 48.5k 1.5k IconSount 8.1k $\pm$ 3.3k
MMSVG-Character 350k 48.9k 1.1k Freepik & generated 28k $\pm$ 7.3k
Table 2: SVG draw-commands. Draw commands used in this work along with their arguments and a visualization are listed. The start-position $(x_{1}$, $y_{1})$ is implicitly defined as the end-position of the preceding command.
Command Arguments Description Visualization
<SOP> $\varnothing$
’Start-of-Path’ token.
M
$x_{2}$, $y_{2}$ Move the cursor to the end-point $(x_{2},y_{2})$ without drawing anything. [Uncaptioned image]
L
$x_{2}$, $y_{2}$ Draw a line to the point $(x_{2},y_{2})$. [Uncaptioned image]
C
(Cubic
Bézier)
$q_{x1}$, $q_{y1}$
$q_{x2}$, $q_{y2}$
$x_{2}$, $y_{2}$
Draw a cubic Bézier curve with control
points $(q_{x1},q_{y1})$, $(q_{x2},q_{y2})$ and end-point
$(x_{2},y_{2})$.
A
(Elliptical
Arc)
$r_{x}$, $r_{y}$
$\varphi$, $f_{A}$, $f_{S}$
$x_{2}$, $y_{2}$
Draw an elliptical arc with radii $r_{x}$ and $r_{y}$
(semi-major and semi-minor axes),
rotated by angle $\varphi$ to the x-axis,
and end-point ($x_{2}$, $y_{2}$). $(x_{2},y_{2})$.
Z (ClosePath) $\varnothing$
Close the path by moving the cursor back
to the path’s starting position $(x_{0},y_{0})$.
F
(Fill) $fill$ Draw the fill attribute of the path. $\varnothing$
<EOS> $\varnothing$
’End-of-SVG’ token.
Data Curation.
We extract SVG samples with a comprehensive deduplication process based on filenames, SVG code, and metadata. We first fit the collected SVGs within a viewbox of $200\times 200$. Then, we employ an off-the-shelf VLM, specifically BLIP-2 [26], to generate captions for the SVGs. Please find more samples from the MMSVG-2M dataset in Fig. 10, and instruction templates in Sec. A.2.
SVG Simplification
is an essential procedure in SVG data cleansing, since the over-complicated XML grammars in the crawled SVG data will lead to ambiguities while representing basic shapes. For example, the Rect command creates a rectangular shape controlled by the starting point, width, and height arguments, like <rect x="80" y="90" width="100" height="100"/>, which can also be represented with four orthogonal lines. Also, the Circle Command creates a circle shape based on a center point and a radius, like <circle cx="50" cy="50" r="50"/>, which can also be approximated with the Bézier curve. Moreover, the Transform attribute in XML grammars applies an affine transformation to an existing shape. For example, <rect x="50" y="5" width="40" height="40" fill="blue" transform="rotate(45)"/> refers to rotating the blue rectangle with $45$ degrees.
To standardize training and evaluation, we simplify all SVG commands with atomic commands as shown in Tab. 2. Inspired by FIGR-8-SVG [11] and IconShop [55], we remove all attributes and simplify each SVG with five basic commands, including “Move To” (M), “Line To” (L), “Cubic Bézier” (C), “Elliptical Arc” (A), “ClosePath” (Z). The introduction of atomic commands further removes the ambiguities, as complex XML grammars can be approximated with the combination of several atomic commands. To efficiently produce a unified and less complex data structure, we utilize picosvg https://github.com/googlefonts/picosvg to remove grammars like “group” and “transform”, and simplify the complex commands to atomic path commands. It is worth noting that atomic path commands are sufficient to represent complex SVGs shown in Fig. 1.
3.2 MMSVG-Bench
As a multi-modal SVG generation evaluation protocol, MMSVG-Bench put emphasis on both the generation quality and how well the model follows multi-modal instructions.
Text-to-SVG requires a model to generate SVGs from text instructions. We measure the visual quality with Frechet Inception Distance (FID) [49], aesthetic appeal with Aesthetic score [42], text-SVG alignment with CLIP score [38], and Human Preference Scores (HPS) [57].
Image-to-SVG evaluates a model’s ability to convert images into SVGs. To quantify the distance between the input and output SVG, we calculate the cosine similarity of DinoV2 features (DinoScore) [34], Structural Similarity Index (SSIM) [53], Learned Perceptual Image Patch Similarity (LPIPS) [66], and Mean Squared Error (MSE).
Character-Reference SVG Generation evaluates a model’s ability to generate novel SVGs while keeping the profile of the characters depicted in the input image. Different from image-to-SVG, the model does not reconstruct, but generates a specific character SVG for the input image (see Fig. 6).
We evaluate the alignment between input character images and generated SVGs by prompting GPT-4o [20] to generate a score ranging from 1 to 10, the higher the better. For evaluation metrics, we also calculate the average token length (# tokens) of a generated SVG sample utilizing the Qwen2.5-VL [1] tokenizer. For computational cost, we compute the average generation time of generating per SVG sample.
4 OmniSVG
To support end-to-end training for multi-modal SVG generation, OmniSVG parameterizes a series of atomic path commands representing an SVG into a sequence before feeding into a pre-trained VLM with multi-modal instructions.
SVG Tokenizer. As illustrated in Sec. 3, our MMSVG-2M dataset simplifies an SVG by removing all attributes and using five basic path commands (see Tab. 2). After the simplification, an SVG script $G$ is represented as the combination of $M$ paths, $G=\{P_{i}\}_{i=1}^{M}$. Here, $P_{i}$ is the $i$-th path containing $N_{i}$ commands, $P_{i}=\{C_{i}^{j}\}_{j=1}^{N_{i}}$, where $C_{i}^{j}$ is the $j$-th command in the $i$-th path. Each command is represented as $C_{i}^{j}=(U_{i}^{j},V_{i}^{j})$, containing both the type identifier $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z}\}$ and the corresponding location argument $V_{i}^{j}$. To generate colored SVG contents, we assign special tokens for hex values to control the “Fill” ($\mathrm{F}$) attribute, distinguishing it from the original SVG commands and coordinates. To this end, we are able to use a total six types of commands $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$ to parameterize a colored SVG parameterization. Specifically, our SVG tokenizer transforms SVG scripts $X_{s}$ into an ordered SVG token sequence within the same representation space as the pre-trained VLM. Following IconShop [55], we flatten the layered structure of the SVG script by concatenating different paths into a single command sequence, where each path begins with the drawing commands followed by point coordinates. Therefore, each SVG sequence could be represented as a flattened sequence. As the generation identifier, we apply special tokens like <SOP> and <EOS> to the two ends of a SVG sequence, identifying the begining and ending of a SVG sequence. We assign special tokens for each command type, i.e. $\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$. To shorten the length of the SVG sequence, we further merge the point coordinates into one token with a mapping function: $<x,y>\rightarrow x\times w+y$, where $w$ is the width of the image. The SVG sequence are then lifted into the same embedding space as the pre-trained VLM with a learnable embedding layer.
Model Architecture. OmniSVG leverages the Qwen2.5-VL architecture, a vision-language models (VLMs), which excels in processing both visual and textual inputs, enabling the generation of precise and compact SVG outputs. OmniSVG is trained for text-to-SVG, image-to-SVG, and character-reference SVG generation task. As depicted in Fig. 2, We first tokenize and embed the interleaved text and image inputs as prefix tokens. Then, we tokenize the SVG scripts into a sequence and concatenate it to the end of the prefix token with the SVG tokenizer. The full sequence is then served as the input of the decoder-only language model.
Training Objective. Similar to LLMs, we train the model to generate new tokens conditioned on the provided prefix tokens with the next-token prediction loss.
$\theta^{*}=\arg\max_{\theta}\prod_{i=1}^{L}P\left(x_{s,i}\mid x_{s,<i},x_{c}\right)$ (1)
Scaling Up. To study the effectiveness of scaling up multimodal SVG generation, we scale up OmniSVG from 3B to 7B parameters. We present training perplexity in Fig. 3, where both models are trained from scratch on 250 billion tokens. We show that, as the size of the model grows, the model achieves a lower validation perplexity, indicating a higher probability of producing the validation data.
(a) Training PPL for our models.
(b) Validation PPL for our models.
Figure 3: Training and Validation Perplexity (PPL) for OmniSVG Models. We train all the models from scratch on 250 billion tokens. We observe that the performance grows with model sizes.
5 Experiments
Table 3: Quantitative Evaluations. Quantitative results between OmniSVG and current state-of-the-art text-to-SVG and image-to-SVG baseline methods. The bold numbers and underlined numbers represents the best and second best performance repectively. Our OmniSVG model demonstrates superior performance compared SOTA SVG generation baselines. StarVector(8B) fails to generate complex SVGs when evaluated in the complex MMSVG-Character dataset.
Evaluation Dataset Methods Time # Tokens Text-to-SVG Image-to-SVG
[…]
Vectorfusion [21] 69s 40k 88.74 0.2798 4.98 0.225
SVGDreamer [59] 260s 100k 75.31 0.2923 5.32 0.256
Chat2SVG [56] 28s 1.7k 70.39 0.3029 5.38 0.256
IconShop [55] 6.8s 3.3k 80.84 0.2359 3.47 0.138
LIVE [33] 170s 18.2k 0.960 0.979 0.034 0.004
DiffVG [27] 21s 19.8k 0.951 0.956 0.056 0.015
GPT-4o [20] 6.9s 0.6k 0.887 0.826 0.179 0.134
StarVector(8B) [41] 45s 3.5k 0.977 0.953 0.048 0.012
OmniSVG(3B) 15s 3.8k 64.83 0.3194 5.49 0.247 0.980 0.954 0.049 0.011
MMSVG-Icon OmniSVG(7B) 23s 4.6k 62.16 0.3278 5.63 0.258 0.984 0.973 0.037 0.008
Vectorfusion [21] 70s 41k 90.03 0.2639 4.67 0.214
SVGDreamer [59] 260s 102k 83.02 0.2795 5.09 0.219
Chat2SVG [56] 35s 2.1k 78.84 0.2891 5.01 0.221
IconShop [55] 7.8s 3.3k 84.04 0.2198 3.21 0.098
LIVE [33] 250s 18.2k 0.959 0.960 0.044 0.011
DiffVG [27] 31s 19.8k 0.929 0.930 0.077 0.021
GPT-4o [20] 8.0s 0.7k 0.801 0.783 0.244 0.199
StarVector(8B) [41] 65s 5.3k 0.807 0.831 0.248 0.181
OmniSVG(3B) 49s 9.7k 70.45 0.3077 5.39 0.245 0.974 0.944 0.069 0.019
MMSVG-Illustration OmniSVG(7B) 63s 10.9k 66.91 0.3164 5.59 0.253 0.988 0.959 0.041 0.013
Vectorfusion [21] 69s 40k 93.03 0.2218 4.08 0.198
SVGDreamer [59] 260s 100k 86.01 0.2423 4.81 0.205
Chat2SVG [56] 42s 1.7k 82.98 0.2499 4.83 0.209
IconShop [55] 6.8s 3.3k 88.41 0.2001 3.13 0.077
LIVE [33] 190s 18.2k 0.911 0.939 0.109 0.038
DiffVG [27] 21s 19.8k 0.876 0.915 0.163 0.051
GPT-4o [20] 8.1s 0.7k 0.712 0.693 0.365 0.349
StarVector(8B) [41]
OmniSVG(3B) 105s 30.8k 72.18 0.2898 5.19 0.213 0.921 0.917 0.049 0.021
MMSVG-Character OmniSVG(7B) 139s 35k 68.02 0.3088 5.33 0.238 0.932 0.941 0.072 0.008
Figure 4: Qualitative comparison with SOTA methods on Text-to-SVG task. We compare the propose method with SOTA Text-to-SVG methods on our evaluation benchmarks, namely Icon, Illustration and Character. The proposed method outperforms existing state-of-the-art approaches in both instruction-following and the aesthetic quality of the generated SVGs.
Figure 5: Qualitative comparison with SOTA methods on Image-to-SVG task. We compare the propose method with SOTA Image-to-SVG methods on our evaluation benchmarks. Despite generating plausible results on simple icon samples, optimization based methods like DiffVG [27] and LIVE [33] tend to output artifacts on complex images. GPT-4o [20] is only able to generate icon-level SVG even given the complex input image. StarVector [41] is able to generate SVG for the input icon image. However, when inputed the illustration or more complex character image, StarVector fails to generate SVGs. Please zoom-in for more details.
To validate the effectiveness of our method, we first introduce the baselines and implementation details (Sec. 5.1). Then, we make quantitative comparisons with prior arts (Sec. 5.2) and ablation studies (Sec. 5.3) to study the effectiveness of our design.
5.1 Baselines
For the text-to-SVG task, we compare our method with language-based (LLM-based) methods, including VectorFusion [21], SVGDreamer [59], Chat2SVG [56] and IconShop [55]. For image-to-SVG task, we compare our method with baseline methods across image vectorization and Multimodal Large Language Modeling approaches, including LIVE [33], DiffVG [27], StarVector [41] and GPT-4o [20] using the official implementations with the hyperparameters proposed by the authors, and apply their pre- and post-processing code as required.
Specifically, for the text-to-SVG task, the optimization-based method SVGDreamer excels in enhancing editability by employing a semantic-driven image vectorization process that effectively separates foreground objects from the background, while failing to handle complex scenes. Another optimization-based work, VectorFusion, stands out for generating SVG-exportable vector graphics without relying on large captioned datasets. However, Vectorfusion is also unable to handle complex scenarios and diverse styles. The significant problem with these optimization-based works is that the optimization time is too long. Generating an SVG usually takes more than ten minutes, which is too expensive. For the LLM-based method, Chat2SVG integrates Large Language Models (LLMs) with image diffusion models to create semantically rich SVG templates. However, Chat2SVG still needs to optimize the output SVG script from LLM, which introduces increased computational complexity and poses challenges during model training. In comparison, IconShop utilizes a transformer-based architecture to autoregressively model SVG path sequences, demonstrating exceptional performance in simplified icon SVGs, which offers effective solutions for text-to-SVG generation. It can only generate black simple Icon SVGs.
For the image-to-SVG task, we compare our method with the image vectorization methods. LIVE allows progressive and efficient generation of SVGs, optimizing closed vector paths under raster image supervision with shape complexity control. However, LIVE needs to optimize for a long time when generating complex SVGs. DiffVG enables end-to-end differentiability in vector graphics rasterization, improving optimization through anti-aliasing and gradient-based methods while also is computationally expensive due to the complexity of the forward-backward rasterization process. Recently, the Multimodal Large Language Model (MLLM) based method StarVector leverages the visual understanding to apply accurate SVG primitive to the LLM architecture, which also can generate SVGs from both text and image inputs. However, it still fails to generate complex SVGs. Since Starvector [41] has not yet opened up its text-to-SVG model weights, our MMSVGBench does not evaluate Starvector’s text-to-SVG capabilities. MMSVG-Bench also evaluates our methods with VLM methods, GPT-4o, to conduct a comprehensive assessment. We compare our method with these baselines on our MMSVG-2M dataset, from simple MMSVG-Icon datset, a bit complex MMSVG-illustration dataset, to the very complex MMSVG-Character dataset. We report the quantitative results in Tab. 3 and the qualitative results in Fig. 4 and Fig. 5.
Figure 6: Generated SVG with Character-Reference (CRef) by OmniSVG. By training on MMSVG-Character with natural character image and SVG pair data, OmniSVG is capable of generating character SVGs through image references.
5.2 Evaluations and Comparisons
5.2.1 Quantitative Evaluations
We report the evaluation metrics of our OmniSVG and baselines on our MMSVG-2M dataset, including the MMSVG-Icon, MMSVG-Illustration and MMSVG-Character in Tab. 3, including both text-to-SVG and image-to-SVG tasks. As shown in Tab. 3, OmniSVG consistently outperforms all other text-to-SVG baselines across all evaluation metrics, indicating lower FID values, higher CLIP scores, higher aesthetic scores and higher HPS scores. For the image-to-SVG tasks, while image vectorization method LIVE achieves higher SSIM, lower LPIPS and lower MSE scores, our OmniSVG can achieve higher DINO scores, indicating the higher semantic features between the input image and the generated SVGs.
5.2.2 Qualitative Evaluations
Text-to-SVG task As shown in Fig. 5, we compare the performance of our method with several baseline approaches using seven distinct text prompt instructions for the text-to-SVG task. For optimization-based methods, such as SVGDreamer [59] and VectorFusion [21], generating an SVG from a single prompt requires significant computation time. These methods rely on iterative optimization processes, which, while effective for refining SVG details, are computationally expensive and time-consuming. In contrast, language-based methods like IconShop [55] and Chat2SVG [56] offer faster generation times, as they leverage pre-trained models to generate SVG templates more efficiently. However, each of these methods comes with notable limitations. IconShop, despite producing complete SVG shapes, is restricted to generating black-and-white graphics, which limits its applicability in scenarios requiring color and visual richness. Chat2SVG, although more flexible, is a two-stage process. In this section, we focus on the first stage of Chat2SVG, which is responsible for generating an initial SVG template. While it offers some flexibility, the generated SVGs often lack the level of detail and semantic consistency observed in the outputs from our method.
Our method, OmniSVG, consistently outperforms all other baselines across various text prompts. The generated SVGs not only retain high fidelity to the input instructions but also incorporate rich color and geometric accuracy. Additionally, OmniSVG demonstrates the ability to handle more complex visual cues, enabling the generation of both simple and intricate SVG designs. This qualitative evaluation highlights the superior performance of our approach in terms of both speed and the quality of the resulting SVGs, making it a more effective and versatile solution compared to existing methods.
Image-to-SVG Task. For the image-to-SVG task, we compare our method with classical image vectorization approaches, including DiffVG [27], LIVE [33], VLM-based methods GPT-4o [20] and StarVector [41]. As shown in Fig. 5, our method outperforms these baselines in both the quality and efficiency of SVG generation.
We present a comprehensive qualitative comparison of our proposed method against state-of-the-art (SOTA) Image-to-SVG techniques, evaluated on our established benchmark datasets. The methods under consideration include optimization-based approaches such as DiffVG [27] and LIVE [33], as well as neural-based methods like GPT-4o [20] and StarVector [41]. While image vectorization methods such as DiffVG and LIVE demonstrate impressive results when applied to simple icon images, they exhibit notable deficiencies when tasked with more complex images. These methods tend to generate visual artifacts, indicating that they struggle with preserving the integrity of intricate visual structures, particularly when the input images feature detailed shapes, gradients, or textures. The GPT-4o model, though capable of generating SVG outputs for complex input images, is constrained to producing only icon-level SVGs. This limitation suggests that GPT-4o is primarily designed to handle simpler iconography and lacks the capacity to fully interpret and convert detailed illustrations into scalable vector graphics, which may be a result of its underlying design or training data constraints. StarVector, another VLM-based method, demonstrates strong performance when tasked with generating SVGs from simple icon images. However, its effectiveness diminishes significantly when dealing with more complex inputs, such as illustrations or intricate character images. In these cases, StarVector fails to generate accurate or meaningful SVG outputs, highlighting its limited generalization ability for more diverse and intricate image types.
While our OmniSVG are able to efficiently convert images into high-quality, editable SVGs, from the input icon images, illustration images to the complex character images, highlighting the superior performance in handling diverse and intricate visual cues, setting it apart from traditional vectorization approaches. We provide more visual results in Fig. 8.
Figure 7: Qualitative study on parametrization. Ablation studies on color parametrization (abbreviated as param.) and coordinate (abbreviated ad coord.) paramterization are conducted.
Figure 8: Illustration of the SVG generation capabilities of OmniSVG.
Character-Reference SVG generation task. As shown in Fig. 6, by training on MMSVG-Character with natural character image and SVG pair data, OmniSVG is capable of generating character SVGs through image references. Moreover, we assess the alignment between input character images and generated SVGs using the third-party VLMs assistant GPT-4o [20]. Our findings reveal an average score of approximately $7$, suggesting that the generated IP SVGs exhibit a well degree of similarity to the input reference images.
5.3 Ablation studies
Table 4: Quantitative study on SVG parameterization. Ablation studies on color parametrization (abbreviated as color param.) and coordinate paramterization (abbreviated as coord param.) are conducted.
Methods Text-to-SVG Image-to-SVG # Tokens
[…]
w/o param. 97.01 0.1537 2.873 0.112 0.702 0.668 0.334 0.232 35k
w/o coordinate param. 87.93 0.2193 3.352 0.156 0.783 0.769 0.295 0.160 29.5k
w/o color param. 76.39 0.2632 4.234 0.201 0.901 0.912 0.098 0.028 13.1k
full param. (Ours) 70.45 0.3077 5.39 0.245 0.974 0.944 0.069 0.019 9.7k
We conduct ablation studies on the SVG parameterization, the scaling up model size and the VLM architecture to validate the effectivenss of our design. Most of our ablation experiments are performed on the OmniSVG-3B model for faster iteration, and we empirically find they work well on the larger OmniSVG-7B.
Effectiveness of the SVG Parameterization. In this experimental study, we present a comprehensive comparison between SVG parameterization approach and traditional non-parameterized methods for SVG representation in large language models. The fundamental distinction lies in how coordinate information is tokenized: traditional methods represent each digit as a separate token (e.g., coordinates $(123,456)$ require six tokens plus delimiters), whereas our parameterization method encodes entire coordinates as single semantic tokens, dramatically improving the model’s ability to process complex path information.
We parameterize both the coordinate and color attributes of the SVG. To validate the effectiveness of our experimental design, we conducted the following ablation experiments: parameterizing color while leaving the coordinates unparameterized, parameterizing coordinates while leaving the color unparameterized, not parameterizing either the coordinates or the color, and parameterizing both the coordinates and the color (our method). The results of these experiments on the MMSVG-Illustration dataset are summarized in the Tab. 4 and Fig. 7.
The results in Tab. 4 indicate that incorporating both coordinate and color parameterizations yields the most favorable outcomes across all evaluated metrics. Notably, this configuration achieves the lowest token length, suggesting that our method efficiently utilizes token space. This efficiency implies that our approach can generate complex SVGs with reduced computational resources, which is advantageous for applications requiring scalability and efficiency. Moreover, the qualitative results in Fig. 7 indicates that the both coordinate and color parameterizations significantly outperforms other methods, particularly as SVG complexity increases. No parameterizations method fail to generate SVGs when given the complex image. These experiments underscore the effectiveness of our full parameterization strategy in balancing performance and resource utilization, making it a compelling choice for SVG generation tasks.
Ablation studies on the model size. To analyze whether training a larger model benefits SVG generation, we evaluate OmniSVG base models with different sizes on the MMSVG-2M dataset in Tab. 5. We evaluate OmniSVG with base models of varying sizes on the MMSVG-2M dataset in Tab. 5. We progressively scale up the model size from the FLAN-T5-base [10] (223M) to Qwen2.5-VL-3B-Instruct (3.7B) and Qwen2.5-VL-7B-Instruct (8.3B). We observe that as the model size grows, the quality of the generated samples gets higher, indicating the effectiveness of scaling up generative auto-regressive training on multi-modal SVG generation tasks.
Table 5: Ablation of the model size. As the model size grows, the generated samples are of higher quality.
Methods Input Size Text-to-SVG Image-to-SVG
FID↓ CLIP↑ Aesthetic↑ HPS↑ DINO↑ SSIM↑ LPIS↓ MSE↓
FLAN-T5-Base[10] Text 223M 89.30 0.177 3.437 0.102
FLAN-T5-Large[10] Text 770M 83.92 0.2303 4.102 0.177
FLAN-T5-xl[10] Text 3B 76.92 0.2789 4.642 0.217
blip2-flan-t5-xl[26] Text/Image 3.94B 75.23 0.2881 4.712 0.201 0.891 0.902 0.099 0.046
OmniSVG(3B) Text/Image 3.7B 70.45 0.3077 5.395 0.245 0.974 0.944 0.069 0.019
OmniSVG(7B) Text/Image 8.3B 66.91 0.3164 5.59 0.253 0.988 0.959 0.041 0.013
Ablation studies on the VLM architecture. To illustrate the effectiveness of the VLM architecture, we conducte an ablation study by replacing the VLM architecture with alternative LLM-based architectures, incorporating image encoders such as CLIP ViT-B/32 [37], VQGAN [14], and the Qwen2.5-VL [1] model used in OmniSVG. These experiments were carried out on the MMSVG-Illustration dataset, with a focus on comparing the performance of different architectures in generating SVG outputs. Among these architectures, Qwen2.5-VL consistently delivered superior results across all evaluation metrics.
The results in Tab. 6 clearly indicate that none of the alternative architectures outperformed Qwen2.5-VL, highlighting the importance of selecting an architecture with a robust image encoder. Additionally, a key advantage of Qwen2.5-VL is its support for multi-image input, which enables the generation of more complex outputs, such as character reference SVGs.
Table 6: Ablation of the VLM architecture. we ablate the VLM architecture by replacing it with the LLM architecture with the image encoder, including the CLIP ViT-B/32 [37], VQGAN [14], and the Qwen2.5-VL [1] we use in OmniSVG. We experiment with training experiments on the MMSVG-Illustration dataset. Qwen2.5-VL architecture gives the best results on all metrics.
Vision Language Text-to-SVG Image-to-SVG
[…]
CLIP Qwen2.5 72.243 0.2876 4.57 0.209 0.881 0.872 0.176 0.062
VQGAN Qwen2.5 76.29 0.2739 4.62 0.198 0.890 0.852 0.183 0.065
Qwen2.5-VL-3B-Instruct 70.45 0.3077 5.39 0.245 0.974 0.944 0.069 0.019
Qwen2.5-VL-7B-Instruct 66.91 0.3164 5.59 0.253 0.988 0.959 0.041 0.013
Table 7: User Study of OmniSVG and baselines. Our OmniSVG achieves the user preference among the baselines.
Method Preference Vividity Alignment
Vectorfusion [21] 40 60 80
SVGDreamer [59] 40 70 80
Chat2SVG [56] 60 60 90
IconShop [55] 80 60 80
GPT-4o [20] 40 60 80
StarVector(8B) [41] 40 80 70
DiffVG [27] 90 80 100
LIVE [33] 90 70 100
OmniSVG [33] 100 90 100
5.4 User study and Implementation Details.
User Study. For user study, we assess the user preference, vividity, and text-SVG/image-SVG alignment of the generated SVGs. In this study, we conduct a user study to evaluate the effectiveness and practicality of the SVGs generated by our method and baselines. Specifically, We randomly sample 5 text descriptions and 3 image prompts from the evaluation dataset and generate corresponding SVGs using our methods and baseline models. Along with the original 8 SVGs from the dataset, this provided a total of 64 SVGs for the user study. In Tab. 7, preference indicates the generated SVG preference, “Vividlity” indicates the sample is more lively, and “Alignment” represents the input text/image instructions more likely to align with the generated SVGs. In this study, we shuffle the SVGs generated by the different models together with the selected examples from our dataset. Participants are only provided with the text descriptions or image prompts and the shuffled SVGs. Tab. 7 also demonstrates that OmniSVG is preferred by volunteers, who perceive the generated SVGs as more vivid and visually aligned with the input text or image. This alignment highlights OmniSVG’s superior ability to produce more accurate and contextually relevant SVG content according to volunteers’ feedback.
Implementation Details. We train our models in bfloat16 with the ZeRO-2 strategy [39] for memory-efficient training. We also adopt the AdamW [32] optimizer with a learning rate decaying from $3\times 10^{-4}$ to $3\times 10^{-6}$ and a weight decay of $0.1$ to train our model. In practice, we load the pre-trained weights from the Qwen2.5-VL [1] model and initialize the SVG embeddings from scratch. Without further specification, we generate SVGs with the top-k and top-p sampling strategy with $k=50$ and $p=0.95$ for diversity.
6 Conclusions
Figure 9: Limitation of OmniSVG on Image-to-SVG Task. OmniSVG can successfully generate vector style images, while fail to fit natural images.
Conclusions. We introduce OmniSVG, a unified framework for multimodal SVG generation that leverages pre-trained Vision-Language Models (VLMs). By parameterizing SVG commands and coordinates as discrete tokens, OmniSVG efficiently decouples structural logic from geometry, addressing issues like ”coordinate hallucination” while maintaining design expressiveness. Our method outperforms existing approaches in both quality and efficiency, offering high-quality, editable SVG across various design domains. Additionally, we proposed MMSVG-2M, a large-scale multimodal dataset with two million annotated SVG assets and a standardized evaluation protocol. Extensive experiments show that OmniSVG surpasses prior SVG generation methods in various conditional generation tasks, highlighting its potential for integration into professional SVG design workflows.
Limitations and Future Work. During inference, OmniSVG generates tens of thousands of tokens for complex samples, which inevitably leads to a considerable generation time. OmniSVG is only bounded by vector style image prompt and fails on natural images as illustrated in Fig. 9. As for future work, recent endeavors on multi-token prediction [15, 2] and KV-cache compression [68, 3] provide a promising way to save the generation cost. Additionally, the auto-regressive nature of OmniSVG also unlocks future opportunities for in-context learning [67, 69, 46], chain-of-thought reasoning [54, 16], and multi-turn interleaved generation [19, 30], thereby providing a more dexterous user control. Moreover, more complicated tasks like stylization [65, 70] is also worth exploring in the future.
Appendix
Appendix A Additional Details of MMSVG-2M dataset
A.1 Samples of MMSVG-2M Dataset
We visualize samples of our MMSVG-2M dataset in Fig. 10. In our MMSVG-2M dataset, 55% of the SVG samples belongs to the MMSVG-Icon, 25% belongs to the MMSVG-Illustration, and the rest 20% belongs to the MMSVG-Character. Among the SVG samples within the MMSVG-Character category, half of them comes from Freepik, while another half is generated by our data creation pipeline. We also collect image-SVG pairs for the character-reference SVG generation tasks during the generation process.
Figure 10: Samples from MMSVG-2M dataset. The proposed MMSVG-2M dataset can be separated into three subset, namely Icon, Illustration and Character. Samples from Icon, Illustration and part of Character subsets are downloaded from Internet. Another part of Character subset is generated by our data creation pipeline, which can provide image and SVG pairs for image prompting task.
A.2 SVG-Image-Text Pairs Construction
Our MMSVG-2M dataset comprises two million SVG samples with the corresponding rasterized images. Following the recent dataset construction pipeline [28], we generate captions on the rasterized images with BLIP-2 [26], thereby providing textual descriptions that enable us to fine-tune our model to follow these instructions. We use CairoSVG [23] for rasterization and remove samples that produced completely white images.
Figure 11: Image prompting dataset creation of MMSVG-2M Character. By utilizing FLUX-Redux and SVG vectorization tools, image prompting data pairs can be generated. We adpot FLUX-Redux downsampling scale with $2,3$ in practice by trading-off the character similarity and complexity of generated SVG.
Annotation. We employ an off-the-shelf VLM, specifically BLIP-2 [26], to generate SVG captions with the prompt below. To reduce hallucinations, we drop the samples with CLIP scores less than 30. We also visualize the distribution annotated keywords of MMSVG-2M dataset in Fig. 12 with word cloud format. And the instruction template for annotation is shown in Tab. 8.
Figure 12: Word cloud visualization of label distribution in the MMSVG-2M dataset. The size of each label corresponds to its frequency of occurrence. The larger the label, the more frequently it appears in the dataset.
Instruction templates. MMSVGBench provides three tasks, including text-to-SVG task, image-to-SVG task and character-reference SVG generation task. Each task needs different instruction templates. For the text and image conditioning SVG generation, we provide the input text or image with VLM architecture. For character-reference SVG generation, we provide the natural charecter reference image and the original image with the VLM architecture. The list of instruction templates for different tasks are shown in Tab. 8.
Employed BLIP2 for SVG Captioning: You are a helpful assistant. Your task is to describe this image in a single sentence, including the object, its color, and its overall arrangement. For example: “Yellow cheers with glasses of alcohol drinks.” / “Heart emojis represent love on Valentine’s Day.” Text-to-SVG: You are a helpful SVG Generation assistant, designed to generate SVG. We provide the text description as input, generate SVG based on the text. Image-to-SVG: You are a helpful SVG Generation assistant, designed to generate SVG. We provide an image as input, generate SVG for this image. Character-Reference SVG Generation: You are a helpful SVG Generation assistant, designed to generate SVG. We provide two image as input, the second image is the character reference image of the first image, generate the character reference SVG based on these two input images.
Table 8: The list of instructions for different tasks, including annotation, text-to-SVG, image-to-SVG and character-reference SVG generation.
A.3 Character-SVG Pairs Construction
As illustrated in the Fig. 1, part of our proposed MMSVG-2M-Character subset is constructed using a generative pipeline. As shown in the pipeline diagram in Fig. 2, we employ a FLUX [24]-based generative model enhanced with a vector-style LoRA to enable the generation of SVG-style data. For image-based conditioning, we adopt FLUX-Redux [25], which injects image features via a SigLIP encoder and projects them into image embeddings. These embeddings are then concatenated with the text tokens as conditioning inputs for FLUX [24]. However, in practice, the original Redux [25] conditioning proves to be overly strong. To address this, we adopt a community-implemented variant of Redux that downsamples the image embeddings in 2D space. As observed in our experiments shown in Fig. 11, a downsampling factor between 2×2$\times$2 × and 3×3$\times$3 × yields the most reasonable SVG-style character references. Finally, we employ VTracer [12] to perform near-instant vectorization of the generated images. To construct the MMSVG-2M-Character subset, we first filter $103k$ character instances from the Danbooru [13] dataset and apply the aforementioned pipeline with motion and expression keywords like previous works [9, 8, 35, 64]. We compare the raw FLUX [24] outputs and their vectorized counterparts, retaining only those samples with PSNR and SSIM scores above a certain threshold as valid data.
Appendix B More details of the baselines
B.1 Text-to-SVG Task
SVGDreamer [59] uses a semantic-driven image vectorization (SIVE) process to separate foreground objects and background, improving editability. The SIVE process utilizes attention-based primitive control and an attention-mask loss function to manipulate individual elements effectively. To address issues in existing text-to-SVG generation methods, the proposed Vectorized Particle-based Score Distillation (VPSD) approach models SVGs as distributions of control points and colors, improving shape, color diversity, and convergence speed.
VectorFusion [21] leverages a text-conditioned diffusion model trained on pixel representations to generate SVG exportable vector graphics without needing large captioned SVG datasets. By optimizing a differentiable vector graphics rasterizer, it distills semantic knowledge from a pretrained diffusion model and uses Score Distillation Sampling to generate an SVG consistent with a caption. Experiments show that VectorFusion improves both quality and fidelity, offering a variety of styles such as pixel art and sketches.
Chat2SVG [56] proposes a hybrid framework that combines the strengths of Large Language Models (LLMs) and image diffusion models for text-to-SVG generation. The approach first uses an LLM to create semantically meaningful SVG templates from basic geometric primitives. A dual-stage optimization pipeline, guided by image diffusion models, refines paths in latent space and adjusts point coordinates to enhance geometric complexity.
IconShop [55] uses a transformer-based architecture to encode path commands and learn to model SVG path sequences autoregressively. It has shown excellent results in simplified icon scenarios and provides a good solution to Text-to-SVG generation by extending the FIGR-8-SVG dataset with captions. We have access to their dataset and original splits and have trained our model on that data using a pre-trained checkpoint (trained on OmniVG dataset). We have extracted the results from IconShop and included them here to compare our method.
LLM4SVG [58] is a framework that leverages Large Language Models (LLMs) to understand and generate Scalable Vector Graphics (SVGs). It employs a structured SVG encoding approach, utilizing learnable semantic tokens to accurately represent SVG components and their properties. This design enables LLMs to produce SVGs that are both semantically aligned with textual descriptions and visually coherent. However, LLM4SVG also has a maximum token length of 2048, limiting its ability to generate highly complex SVGs that require longer sequences.
B.2 Image-to-SVG Task
LIVE (Layer-wise Image Vectorization) [33] is a method for progressively generating SVGs that closely fit a given raster image by recursively adding and optimizing closed vector paths. Using a differentiable renderer (based on DiffVG [27]), LIVE enables direct optimization of paths under raster image supervision while controlling shape complexity by adjusting the number of path segments. It introduces component-wise path initialization, identifying key visual components to ensure efficient topology extraction and minimize redundant shapes.
DiffVG [27] is a landmark in vector graphics research, pioneering deep learning-based methods with the first differentiable vector graphics rasterization pipeline. By leveraging a combination of anti-aliasing techniques and gradient-based optimization, DiffVG ensures differentiability. Unlike methods relying on non-differentiable curve-to-mesh conversions, DiffVG employs a forward-backward rasterization process, where the forward pass generates antialiased images and the backward pass computes gradients with respect to vector graphic parameters.
StarVector [41] works directly in the SVG code space, leveraging visual understanding to apply accurate SVG primitives. StarVector employs a transformer-based architecture that integrates an image encoder with a language model, enabling it to process visual inputs and produce precise SVG code. StarVector effectively handles diverse SVG types, including icons, logos, and complex diagrams, demonstrating robust generalization across various vectorization tasks. However, with a 16k token context window, StarVector may struggle to process highly complex SVGs that require longer sequences.